OpenStudy (anonymous):

Solve by factoring 6t^2 + t – 5= 0

OpenStudy (nurali):

6t^2 + t – 5= 0 (6t-5)(t+1)=0 6t-5=0 or t+1=0 t=5/6 or t=-1